OpenStudy (anonymous):

how do you do this question? Sarah pulls an object of 23kg with a force of 35N. Shelley pulls the same object at the same time in the opposite direction with a force of 53N. What is the magnitude of the resulting acceleration if the object is on a frictionless surface?

OpenStudy (anonymous):

F=ma is the formula to find acceleration. a=F/m = 35/23= 1.52 m/s2 for 2nd force a=f/m = 53/23 = 2.30 m/s2 both acceleration have opposet direction so take their difference of both accelerations... 2.30 -1.52 =0.78 m/s2

OpenStudy (gebooors):

It is possible first to calculate net force. In this case it is 53 N - 35 N = 18 N. Direction is towards Shelley. Then a = F/m